# Dirichlet process mixtures of block $g$ priors for model selection and prediction in linear models

### v1

@@ 6.2 Model selection, estimation and prediction performance
Taken together, the results in Table 1 and Figure 4 highlight (a) the benefits of using differential shrinkage priors in the context of model selection and (b) the risks associated with the use of fixed rather than data-driven blocks in the development of model selection priors: there are very small gains in efficiency (if any) that are possible if the blocking structure are correctly identified, and very big potential losses if they are not.
(a) Dataset 31
(b) Dataset 63
Figure 6: Joint and marginal posterior distributions for $p_{\gamma}$ and $K_{\boldsymbol{\gamma}}$ in two representative examples of our simulated datasets in the $p=250$ and $\eta=0$ scenario.
Next, Figure 6, which shows joint and marginal posterior distributions for $p_{\gamma}$ and $K_{\boldsymbol{\gamma}}$ in two representative datasets, provides insights into the adaptive properties of the DP block-$g$ prior. Note that, in both cases, the posterior distribution puts most of its mass in at least two clusters of coefficients, with a (marginal) mode of 3 and maximum of 9 in the Dataset 63, and mode of 6 and a maximum of 13 in Dataset 31. These results are in agreement with the theory developed in Theorem 4.3.
@@
Finally, we present in Figure 5 the prediction mean square error (MSE) for each of the procedures. To obtain these prediction MSEs, each dataset was augmented with a test set of 500 additional observations. Furthermore, in order to simply interpretation, we compute the relative MSE with respect to that under the $g$-prior for each dataset. Hence, values less than 1 correspond to methods with smaller (better) prediction MSE. Note that, with the exception of ALasso, all procedures consistently outperform $g$-priors, again illustrating the the advantages of differential shrinkage in Bayesian settings. Additionally, note that Horseshoe has worse performance than the other Bayesian procedures, particularly in “large $p$” scenarios. While perhaps surprising at first sight, this observation is consistent with results in Lee et al. (2020). The remaining five procedures perform similarly. However, Som et al. $(K=3)$ seems to perform slightly better than Som et al. $(K=2)$ in this evaluation, specially when $\eta=0.5$. This is the opposite of what we observed when evaluating model selection performance. Similarly, DP block-$g$ also seems to perform slightly better than GL-$g$, although the differences are somewhat difficult to see because of the logarithmic scale. Similar patterns hold for the mean squared errors of the point estimators of the coefficients (please see Section G of the supplementary materials)
@@ 7 The ozone dataset
We further investigate the performance of DP mixtures of block $g$ prior using the ozone dataset introduced in Breiman & Friedman (1985) and later analyzed in Casella & Moreno (2006) and Liang et al. (2008), among others. The dataset consists of daily measurements of the maximum ozone concentration near Los Angeles and eight meteorological variables. We consider regression models that might include all eight of these variables along with all possible interactions and squares, leading to up to 44 possible predictors.
@@
Finally, Figure 9 presents boxplots of the predictive mean squared error (MSE) and 95% median intervals scores (MIS, Gneiting & Raftery, 2007) for a crossvalidation exercise in which 20 random 80-20 splits of the data were used train and then test prediction accuracy. For a variable $z$, the $\alpha\times 100\%$ IS is given by $IS_{\alpha}(l,u,z)=(u-l)+\frac{2}{1-\alpha}(l-z)\mathbbm{1}\{z<l\}+\frac{2}{1-\alpha}(z-u)\mathbbm{1}\{u<z\}$, where $l$ and $u$ denote the upper and lower bounds of the $\alpha\times 100\%$ posterior intervals of $z$. The first term in this expression rewards narrow predictive intervals, while the second rewards accurate coverage. We do not report the MIS for ALasso because the implementation in the glmnet package does not provide a measure of predictive uncertainty off the shelf. Generally speaking, Horseshoe, HSM and ALasso seem to have a slightly better predictive performance than the Bayesian procedures, particularly when it comes to point prediction. However, the differences between the various procedures are small.
@@ 8 Discussion
We introduced Dirichlet process mixtures of block $g$ priors as a parsimonious, data driven approach to model selection and prediction in linear models that is free from the so-called conditional Lindley “paradox” and that provides a bridge between two strands of the literature (model selection priors and continuous shrinkage priors) that have often been treated as distinct. We also demonstrate that MCMC implementations of models based on DP mixtures block $g$ prior are possible that have the same computational complexity as that of standard $g$-priors and require minimal parameter tuning.
@@
While this paper focuses on linear regression models, DP mixtures of block $g$ priors can be extended to generalized linear models, and perhaps even non-linear regression, by setting $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}$ to be an appropriate information matrix, e.g., see Bové & Held (2011), Li & Clyde (2018) and Porwal & Rodríguez (2023). The approach introduced here can also be used to generalize the class of priors introduced in Carvalho & Scott (2009), leading to a new class of (mixtures of) hyper-inverse Wishart block $g$ priors for model selection in Gaussian graphical models.

### v2

@@ 6.2 Model selection, estimation and prediction performance
Taken together, the results in Table 1 and Figure 3 highlight (a) the benefits of using differential shrinkage priors in the context of model selection, (b) the need to either properly center the distribution of the shrinkage coefficients or, alternatively, allowing enough model flexibility to learn its center from its data, and (c) the risks associated with the use of fixed rather than data-driven blocks in the development of model selection priors.
@@
Finally, we present in Figure 4 the prediction mean square error (MSE) for each of the procedures. To obtain these prediction MSEs, each dataset was augmented with a test set of 500 additional observations. Furthermore, in order to simply interpretation, we normalized all MSEs with respect to that under the $g$-prior for each dataset and graph the resulting ratios in a logarithmic scale. This means that values less than 1 correspond to methods with smaller (better) prediction MSE. Note that, with the exception of ALasso, all procedures consistently outperform $g$-priors in “large pp” regimes, again illustrating the the advantages of differential shrinkage in Bayesian settings. Note that Horseshoe has worse performance than the other Bayesian procedures, particularly in “large pp” scenarios. While perhaps surprising at first sight, this observation is consistent with the results of Lee et al. (2020). Som et al. $(K=3)$ seems to perform slightly better than Som et al. $(K=2)$ in this evaluation, specially when $\eta=0.9$, which is the opposite of what we observed when evaluating model selection performance. Similar to model selection performance, procedures based of the GL-$g$ ($\tau^{2}=1$) prior also have worse performance than the remaining GL-$g$ and DP block-$g$ procedures, all of which perform quite similarly in terms of predictions. Additional results related to simulation study, including results for $\eta=0.5$, examples of posterior distributions over $(p_{\boldsymbol{\gamma}},K_{\boldsymbol{\gamma}})$, and a sensitivity analysis for priors on the concentration parameter $\alpha$ and model space $\boldsymbol{\gamma}$ can be seen in the supplementary materials.
@@ 7 The ozone dataset
We further investigate the performance of DP mixtures of block $g$ prior using the ozone dataset introduced in Breiman and Friedman (1985) and later analyzed in Casella and Moreno (2006) and Liang et al. (2008), among others. The dataset consists of daily measurements of the maximum ozone concentration near Los Angeles and eight meteorological variables. We consider regression models that might include all eight of these variables along with all possible interactions and squares, leading to up to 44 possible predictors.
@@
Finally, Figure 7 presents boxplots of the predictive mean squared error (MSE) and 95% median intervals scores (MIS, Gneiting and Raftery, 2007) for a crossvalidation exercise in which 20 random 80-20 splits of the data were used to train and then test prediction accuracy. For a variable $z$, the $\alpha\times 100\%$ IS is given by $IS_{\alpha}(l,u,z)=(u-l)+\frac{2}{1-\alpha}(l-z)\mathbbm{1}\{z<l\}+\frac{2}{1-\alpha}(z-u)\mathbbm{1}\{u<z\}$, where $l$ and $u$ denote the upper and lower bounds of the $\alpha\times 100\%$ posterior intervals of $z$. The first term in this expression rewards narrow predictive intervals, while the second rewards accurate coverage. We do not report the MIS for ALasso because the implementation in the glmnet package does not provide a measure of predictive uncertainty off the shelf. Generally speaking, Horseshoe, HSM and ALasso seem to have a slightly better predictive performance than procedures based on standard $g$-priors, GL-$g$ priors and DP block $g$ priors, particularly for point prediction. However, the differences are small.
@@ 8 Discussion
We introduced novel class of priors that that enable a parsimonious, data driven approach to model selection and prediction in linear models that is free from the so-called conditional Lindley “paradox” and that provides a bridge between two strands of the literature (model selection priors and continuous shrinkage priors) that have often been treated as distinct. The use of a Dirichlet process prior for the distribution of the local shrinkage coefficients enable us to interpret our model in terms of blocks of covariates that receive the same level of shrinkage. However, the method is not specially reliant to the clustering properties of Dirichlet process. In particular, it does not require that we are able to consistently estimate the “true” partition of the coefficients, only that the partitions that receive high probability a posteriori be those that do not mix both “small” and “large” coefficients.
@@
While this paper focuses on linear regression models, DP mixtures of block $g$ priors can be extended to generalized linear models, and perhaps even non-linear regression, by setting $\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}$ to be an appropriate information matrix, e.g., see Bové and Held (2011), Li and Clyde (2018) and Porwal and Rodríguez (2023). The approach introduced here can also be used to generalize the class of priors introduced in Carvalho and Scott (2009), leading to a new class of (mixtures of) hyper-inverse Wishart block $g$ priors for model selection in Gaussian graphical models.
Recently, Lee et al. (2024) demonstrated that, in the context of continuous shrinkage priors, the behavior of the tail of the distribution of the local shrinkage parameters plays an important role in the performance of the prior, and that different types of tail behavior might be necessary in sparse vs. ultra-sparse settings. The use of a non-parametric specification for the distribution of $g_{j}\mid\boldsymbol{\gamma}$ through a Dirichlet process prior enables such tail adaptability in the context of model selection priors.